# Data-driven Error Estimation: Upper Bounding Multiple Errors without Class Complexity as Input

### v3

@@ 3.3 Supervised Learning: Excess Risk Estimation and Constructing Set of Potentially Optimal Models
We consider the supervised learning setting – described by a distribution $\mathcal{D}$ over $\mathcal{X}\times\mathcal{Y}$, a model class $\mathcal{G}\subseteq\mathcal{X}\rightarrow\mathcal{Y}$, a loss function $l:\mathcal{X}\times\mathcal{Y}\times\mathcal{G}\rightarrow[0,M]$ (for some $M>0$), and two independent datasets $S_{\text{def}}$ and $S_{\text{err}}$ each containing $n$ i.i.d. samples from the distribution $\mathcal{D}$. For any model $g$, the risk is defined as $\mathcal{R}(g):=\mathop{\mathbb{E}}_{D}[l(X,Y,g)]$. The optimal model $g^{*}\in\mathcal{G}$ is then a solution to $\min_{g\in\mathcal{G}}\mathcal{R}(g)$ and the excess risk is defined as $\Delta(g):=\mathcal{R}(g)-\mathcal{R}(g^{*})$. Excess risk ($\Delta(g)$) can be interpreted as a measure of room for improvement over the model $g$.
Suppose we are motivated by the need to determine if we should collect more data for learning from the class $\mathcal{G}$. In this case, for some estimated model $g_{\text{def}}\in\mathcal{G}$ (which may depend on $S_{\text{def}}$ but not on $S_{\text{err}}$), we would like to estimate an upper bound on excess risk ($\Delta(g_{\text{def}})$). If the upper bound on excess risk is large, then collecting more data should allow us to learn a better model from the class $\mathcal{G}$, or at least help us decrease uncertainty for model quality. In general, bounding excess risk is a challenging problem, with many existing approaches relying on naive methods or complex theoretical tools. Using the localization technique introduced in Section 2.3, we construct excess risk bounds and are thus also be able to construct a set of potentially optimal models – that is, a set which will contain $g^{*}$ with high-probability. [^16] Constructing a set of potentially optimal models is the primary statistical step when constructing confidence intervals for UCB (upper confidence bound) based contextual bandit/RL algorithms.
We formulate the problem of upper bounding the excess risk ($\Delta(g_{\text{def}})$) as an error estimation problem. For each $g\in\mathcal{G}$, we define our estimand as $\theta_{g}:=\mathcal{R}(g_{\text{def}})-\mathcal{R}(g)=\mathop{\mathbb{E}}_{D}[l(X,Y,g_{\text{def}})-l(X,Y,g)]$. That is, our estimand $\theta_{g}$ measures the sub-optimality of $g_{\text{def}}$, when compared against a candidate model $g$. Therefore, the excess risk $\Delta(g_{\text{def}})$ is equal to $\theta_{g^{*}}$. Now, our estimate is defined as $\hat{\theta}_{g}:=(1/n)\sum_{(x,y)\in S_{\text{def}}}(l(x,y,g_{\text{def}})-l(x,y,g))$ and the corresponding error is $e_{g}:=\theta_{g}-\hat{\theta}_{g}$. We define a corresponding estimate from $S_{\text{err}}$, $\hat{\theta}_{g,\text{err}}$ as outlined in Table 3.
@@
Table 3: Error Estimation Quantities for Excess Risk Estimation
We now construct $\hat{u}$. From Hoeffding’s inequality, we have
$\hat{u}(S_{\text{err}},g,\delta):=\hat{\theta}_{g,\text{err}}-\hat{\theta}_{g}+M\sqrt{\frac{\log(1/\delta)}{2n}}$
satisfies 2.1. Now, defining $\hat{\xi}_{0}(S_{\text{err}},\delta):=\max_{g\in\mathcal{G}}\hat{u}(S_{\text{err}},g,\delta)$, we have by Theorem 2.2, (14) holds with probability at least $1-\delta$.
@@
$\hat{\xi}_{k}(S_{\text{err}},\delta)$ $:=\max_{g\in\mathcal{G}_{k}}\hat{u}(S_{\text{err}},g,\delta);$ (15)
$\mathcal{G}_{k}$ :=$:=\bigg{\{}g\in\mathcal{G}_{k-1}\bigg{|}-\hat{\theta}_{g,\text{def}}\leq\hat{\xi}_{k-1}(S_{\text{err}},\delta)\bigg{\}};$
For any $x\in\mathcal{X}$, this set of potentially optimal models ($\mathcal{G}_{k}$) allows us to construct a high-probability confidence set for $g^{*}(x)$, given by $\{g(x)|g\in\cap_{k\in\mathbb{N}}\mathcal{G}_{k}\}$. [^17] For some $K$, one would observe $\hat{\xi}_{K+1}=\hat{\xi}_{K}$, and stop the localization procedure. Such confidence sets can be helpful for a variety of decision-making tasks. We summarize our results in Corollary 3.5. It is important to note that unlike popular theoretical bounds on excess risk (e.g., Koltchinskii, 2011), none of our results required the complexity of class $\mathcal{G}$ as an input.
@@
Consider the supervised learning setting described in Section 3.3. With probability at least $1-\delta$, we have (16) hold for all $k\in\mathbb{N}$.
$\forall g\in\mathcal{G},\;\;e_{g}\leq\hat{\xi}_{0}(S_{\text{err}},\delta);\;\;\;\;\;g^{*}\in\mathcal{G}_{k};\;\;\;$ (16)
$\forall x\in\mathcal{X},\;\;g^{*}(x)\in\{g(x)|g\in\mathcal{G}_{k}\};$
$\Delta(g_{\text{def}})\leq\max_{g\in\mathcal{G}_{k}}\hat{\theta}_{g}+\hat{\xi}_{k}(S_{\text{err}},\delta).$
@@
Recall $\hat{\theta}_{g}:=(1/n)\sum_{(x,y)\in S_{\text{def}}}(l(x,y,g_{\text{def}})-l(x,y,g))$, and so a reasonable choice of $g_{\text{def}}$ would be the minimizer $g_{\text{def}}=\operatorname*{arg\,min}_{g\in\mathcal{G}}\mathbb{E}_{S_{\text{def}}}[l(X,Y,g)]$. In this case we would have $\hat{\theta}_{g}\leq 0$ and we could reduce our excess risk bound to $\Delta(g_{\text{def}})\leq\min_{k}\hat{\xi}_{k}(S_{\text{err}},\delta)$.
Figure 2 illustrates the performance of our error estimation bound on excess risk compared to the empirical risk of $g_{\text{def}}$ on $S_{\text{err}}$ for increasing noise levels. Empirical risk is an estimate of true risk, which serves as valid yet naive bound on the excess risk. This is because the true risk is the sum of the excess risk and the irreducible noise of the dataset. Our error estimation approach is designed to cancel out irreducible noise via the construction $\hat{u}$, as the performance of $g_{\text{def}}$ is compared to other models in $\mathcal{G}$. As the noise levels increase, our (probabilistic) bound shows increased variance, but remains closer on average to the true excess risk, demonstrating its ability to distinguish between excess risk and irreducible noise. See Section B.1 for details regarding the simulation.
Figure 2: Comparison of excess risk bounds over a class of Neural Networks. We plot the empirical risk of $g_{\text{def}}$ on $S_{\text{err}}$ (blue), our (probabilistic) error estimation bound (orange) and the true excess risk (green) against increasing noise levels.

### v4

@@ 3.3 Supervised Learning: Excess Risk Estimation and Constructing Set of Potentially Optimal Models
We consider supervised learning with data $(X,Y)\sim\mathcal{D}$ over $\mathcal{X}\times\mathcal{Y}$, a model class $\mathcal{G}\subseteq\mathcal{X}\to\mathcal{Y}$, a bounded loss $l:\mathcal{X}\times\mathcal{Y}\times\mathcal{G}\to[0,M]$, and two independent datasets $S_{\text{def}}$ and $S_{\text{err}}$ of $n$ i.i.d. samples each from $\mathcal{D}$. The risk of a model $g$ is $\mathcal{R}(g):=\mathop{\mathbb{E}}_{\mathcal{D}}[l(X,Y,g)]$. Let $g^{*}=\arg\min_{g\in\mathcal{G}}\mathcal{R}(g)$ denote the best-in-class model; the excess risk $\Delta(g):=\mathcal{R}(g)-\mathcal{R}(g^{*})$ quantifies how suboptimal $g$ is.
Given a current model $g_{\text{def}}$, an upper bound on the excess risk ($\Delta(g_{\text{def}})$) would help determine if we should collect more data for learning from the class $\mathcal{G}$. A large bound implies room for improvement, while a small bound indicates the model is already close to optimal. Bounding excess risk typically requires complex theoretical tools. Using the localization technique introduced in Section 2.3, we construct excess risk bounds and thus define a set of potentially optimal models – that is, a set which will contain $g^{*}$ with high-probability. [^12] Constructing a set of potentially optimal models is the primary statistical step when constructing confidence intervals for UCB (upper confidence bound) based contextual bandit/RL algorithms.
We formulate the problem of upper bounding the excess risk ($\Delta(g_{\text{def}})$) as an error estimation problem. For $g\in\mathcal{G}$, let $\theta_{g}:=\mathcal{R}(g_{\text{def}})-\mathcal{R}(g)$ be a measure of the sub-optimality of $g_{\text{def}}$, when compared against a candidate model $g$. Therefore, the excess risk $\Delta(g_{\text{def}})$ is equal to $\theta_{g^{*}}$. We define estimates of $\theta_{g}$ from $S_{\text{def}}$ and $S_{\text{err}}$, as outlined in Table 3.
Table 3: Error Estimation Quantities for Excess Risk Estimation
@@
Now we use Hoeffding’s inequality, to construct a $\hat{u}$ which satisfies 2.1.
$\hat{u}(S_{\text{err}},g,\delta):=\hat{\theta}_{g,\text{err}}-\hat{\theta}_{g}+2M\sqrt{\frac{\log(1/\delta)}{2n}}$
Defining $\hat{\xi}_{0}(S_{\text{err}},\delta):=\max_{g\in\mathcal{G}}\hat{u}(S_{\text{err}},g,\delta)$, by Theorem 2.2, (12) holds with probability at least $1-\delta$.
@@
$\hat{\xi}_{k}(S_{\text{err}},\delta)$ $=\max_{g\in\mathcal{G}_{k}}\hat{u}(S_{\text{err}},g,\delta);$ (13)
$\mathcal{G}_{k}$ $=\bigg\{g\in\mathcal{G}_{k-1}\bigg|-\hat{\theta}_{g}\leq\hat{\xi}_{k-1}(S_{\text{err}},\delta)\bigg\};$
For any $x\in\mathcal{X}$, this set of potentially optimal models ($\mathcal{G}_{k}$) allows us to construct a high-probability confidence set for $g^{*}(x)$, given by $\{g(x)|g\in\cap_{k\in\mathbb{N}}\mathcal{G}_{k}\}$. [^13] For some $K$, one would observe $\hat{\xi}_{K+1}=\hat{\xi}_{K}$, and stop the localization procedure. We summarize our results in Corollary 3.5. It is important to note that unlike popular theoretical bounds on excess risk (e.g., Koltchinskii, 2011), none of our results required an a priori analytic upper bound on class complexity of $\mathcal{G}$ as an explicit input.
(a) 10 dimensional linear regression model class, simulated datataset. Error estimated excess risk bounds (green) compared to VC theory excess risk bounds (red) in the linear case where VC theory is tight. True excess risk is shown in blue.
(b) Neural networks with the structure [input dim - 128 - 64 - 32 - output dim] and ReLU activation. Error estimated excess risk bounds (orange) compared to the naive empirical risk (blue). True excess risk is shown in green.
Figure 2: Performance of error-estimated excess risk bounds in different settings. See Appendix C for simulation details. Figure 2(a) is based on synthetic data, Figure 2(b) uses the Iris dataset (Fisher, R. A., 1936), downloaded from UCI.
@@
Consider the supervised learning setting described in Section 3.3. With probability at least $1-\delta$, we have $\forall g\in\mathcal{G},e_{g}\leq\hat{\xi}_{0}(S_{\text{err}},\delta);$ and for all $k\in\mathbb{N}$, (14) hold.
$g^{*}\in\mathcal{G}_{k};$ $\Delta(g_{\text{def}})\leq\max_{g\in\mathcal{G}_{k}}\hat{\theta}_{g}+\hat{\xi}_{k}(S_{\text{err}},\delta).$ (14)
@@
Note that a reasonable choice of $g_{\text{def}}$ would be the minimizer $g_{\text{def}}=\operatorname*{arg\,min}_{g\in\mathcal{G}}\mathbb{E}_{S_{\text{def}}}[l(X,Y,g)]$. Then $\hat{\theta}_{g}\leq 0$ and we could reduce our excess risk bound to $\Delta(g_{\text{def}})\leq\min_{k}\hat{\xi}_{k}(S_{\text{err}},\delta)$.
Figure 2 highlights two regimes: (a) for linear models our bound coincides with the tight VC-theory curve, confirming no loss versus theory; (b) for neural nets, as noise is added, our method still hugs the true excess risk (albeit with increased variance) while the naïve empirical-risk gap blows up. This demonstrates our method’s ability to distinguish between excess risk and irreducible noise.
4 Applications to Contextual Bandits
In this section, we focus on the stochastic contextual bandit setting, further developed in Appendix H. Contextual bandit algorithms require uncertainty estimates to effectively balance exploration and exploitation. In each round $t=1,\dots,T$, the learner observes a context $x_{t}\in\mathcal{X}$, chooses an action $a_{t}\in\mathcal{A}$, and observes a reward $r_{t}(a_{t})$. The goal is to compete with the best policy in a class $\Pi\subseteq(\mathcal{X}\to\mathcal{A})$:
$\pi^{*}\in\operatorname*{arg\,max}_{\pi\in\Pi}\mathop{\mathbb{E}}_{(x,r)}[r(\pi(x))].$
Computationally tractable and theoretically sound bandit algorithms typically obtain exploration bonuses from error bounds on a reward/value model class $\mathcal{F}$ (e.g., $f(x,a)\approx\mathop{\mathbb{E}}[r(a)\mid x]$). These bounds typically require a complexity measure of $\mathcal{F}$ (VC/Rademacher/covering numbers) as an explicit input to the algorithm (Foster and Rakhlin, 2023; Agarwal et al., 2014; Foster et al., 2020). For many model classes $\mathcal{F}$ (e.g., deep nets, tree-based policies), theoretically deriving complexity bounds is hard and often conservative.
We show error estimation provides a drop-in, data-driven substitute: it yields the needed confidence widths without requiring the complexity of $\mathcal{F}$ as input. We first plug it into a simple inverse gap weighting (IGW) method in the linear setting (where tight theory exists under realizability assumptions141414i.e. the assumption that the true reward model $f\in\mathcal{F}$.). We then (Section 4.2) outline a theoretically sound and modular pipeline where reward models can be estimated via any blackbox oracle [^15] This oracle may not be related to the policy class of interest, for example the oracle could select the best reward model from multiple classes via cross validation. and error estimation is used for constructing two forms of uncertainty sets (conformal arm sets and arm elimination sets) to quantify the error of the reward model in terms of evaluating the policy class of interest without assuming realizability. Details and guarantees for this pipeline are provided in Appendix H.
4.1 Improving FALCON (IGW) via error estimation
Figure 3: Linear contextual bandit: FALCON with standard linear uncertainty vs. FALCON with error-estimated uncertainty. The top plot shows excess risk bounds using theoretical bounds vs. error estimation. The bottom plot shows performance of algorithms based on said excess risk bounds. See Appendix C for details.
Many contextual bandit algorithms follow the same template: fit a reward model from past data, form uncertainty estimates, and construct a policy that balances exploration vs exploitation. A representative example is FALCON (Simchi-Levi and Xu, 2020), which fits a regressor $\hat{f}_{t}$ for $\mathop{\mathbb{E}}[r\mid x,a]$ and quantifies uncertainty via a theoretical excess risk bound over the underlying reward model class – as the excess risk bound shrinks, FALCON samples arms with higher estimated reward more often.
By using data from earlier epochs as $S_{\text{def}}$ and data from the last epoch as $S_{\text{err}}$, we compute a high-probability upper bound $\hat{\xi}$ on the excess risk of the current estimated reward model. As this bound is instance-dependent, it tends to be tighter, thus using it in FALCON results in lower cumulative regret (Figure 3). We illustrate this in a linear setting, where standard theory is already well understood; we expect the benefit to be larger for richer model classes where complexity bounds are often looser.
4.2 A Modular Contextual Bandit Pipeline
[…]
In addition to the modularity and not requiring mathematically derived complexity bounds as an algorithmic input, a key benefit is that the pipeline does not require realizability: we do not assume that the true conditional expected reward model ($\mathop{\mathbb{E}}[r(a)\mid x]$) belongs to our reward estimation class $\mathcal{F}$. Instead, error estimation is used to evaluate how well a given reward model estimates differences in policies within the class of interest [^17] Similar to the measure of error used in misspecification tests by (Krishnamurthy et al., 2024). [^18] Further, since we only evaluate if our reward model can differentiate between policies in the class of interest, our reward models can be estimated via heterogeneous treatment effect estimators and enjoy the corresponding statistical benefits (see, e.g. (Carranza et al., 2023))..
Data from past epochs $S$ $S_{\text{def}}$ $S_{\text{err}}$ black-box oracles $\hat{R},\hat{\pi},\hat{f}$ error estimation $\hat{U}_{\text{arm elimination}}$    and    $\hat{U}_{\text{conformal arm sets}}$ sets for exploration elimination $\hat{g}(x)$    +    conformal $C(x)$ exploration policy $p_{\tau+1}$
Figure 4: Bandit pipeline: error estimation calibrates uncertainty used for both elimination and conformal arm sets, allowing for a modular approach with no realizability assumption required. See Appendix H for algorithm details.